# Machine-learning inference of stellar properties using integrated photometric and spectroscopic data

### v1

@@ Abstract
Stellar astrophysics relies on diverse observational modalities—primarily photometric light curves and spectroscopic data—from which fundamental stellar properties are inferred. While machine learning (ML) has advanced analysis within individual modalities, the complementary information encoded across modalities remains largely underexploited. We present DESA (Dual Embedding model for Stellar Astrophysics), a novel multi-modal foundation model that integrates light curves and spectra to learn a unified, physically meaningful latent space for stars. DESA first trains separate modality-specific encoders using a hybrid supervised/self-supervised scheme, and then aligns them through DualFormer, a transformer-based cross-modal integration module tailored for astrophysical data. DualFormer combines cross- and self-attention, a novel dual-projection alignment loss, and a projection-space eigendecomposition that yields physically structured embeddings.
We demonstrate that DESA significantly outperforms leading unimodal and self-supervised baselines across a range of tasks. In zero- and few-shot settings, DESA’s learned representations recover stellar color–magnitude and Hertzsprung–Russell diagrams with high fidelity ($R^{2}=0.92$ for photometric regressions). In full fine-tuning, DESA achieves state-of-the-art accuracy for binary star detection (AUC = $0.99$, AP = $1.00$) and stellar age prediction (RMSE = $0.94$ Gyr). As a compelling case, DESA naturally separates synchronized binaries from young stars—two populations with nearly identical light curves—purely from their embedded positions in UMAP space, without requiring external kinematic or luminosity information. DESA thus offers a powerful new framework for multimodal, data-driven stellar population analysis, enabling both accurate prediction and novel discovery.
@@ 1 Introduction
Understanding the fundamental properties of stars is central to astrophysics, enabling insights into stellar evolution, galactic structure, and the conditions for planet formation. Traditionally, this effort was done using an analysis of different stellar measurements, such as stellar light curves (photometry) and stellar spectra (spectroscopy). While classical spectra analysis is usually used to predict stellar parameters related to absorption and emission lines such as $T_{\mathrm{eff}},\log g,v\sin i$ and metallicity $[\mathrm{Fe}/\mathrm{H}]$. (García Pérez et al., 2016; Wu et al., 2014), classical light curves analysis usually uses spots modulation to find periodicity signatures (Reinhold et al., 2013; McQuillan et al., 2014; Santos et al., 2019; Lu et al., 2020; Santos et al., 2021; Reinhold et al., 2023; Hattori et al., 2025) and magnetic activity (Mathur et al., 2014; Santos et al., 2024). The revolution of deep learning models also affected stellar astrophysics, with various works using data-driven models that learns to predict or classify stellar parameters from observations and simulations. For example, Blancato et al. (2020), Claytor et al. (2024),Kamai & Perets (2025a), and Claytor & Tayar (2025) used deep learning models to predict rotational period from light curves. Pan et al. (2024b) and Zuo et al. (2025) utilized transformer-based models to predict $\log g$. The use of machine learning in spectral analysis dates back to the pre-deep learning era, with the work of Bailer-Jones (2000). Since then, many works have been done combining deep learning and stellar spectra, with Leung & Bovy (2019), Bai et al. (2020), Olney et al. (2020), Leung & Bovy (2023), Li & Lin (2023), and Koblischke & Bovy (2024) as some examples. For a review on machine learning in astronomy, please refer to Li et al. (2025).
@@
Although the richness of works at the intersection between machine learning and stellar astrophysics, most of them use a single modality, and we refer to them as unimodal models. Multi-modality models, on the contrary, try to combine the information from different modalities of the same object. This approach showed great success in NLP and vision with works like CLIP (Radford et al., 2021) and its variants. Multimodality is of great importance in astrophysics – for example, photometry and spectroscopy reveal partial and very different stellar information; thus, accurate estimation of some stellar properties requires information from both modalities. Despite its importance, there are few works that utilize multi-modal models in astrophysics. AstroCLIP (Parker et al., 2024) is one of such few attempts. They authors used galaxy images and spectra to train a CLIP-like model that aligns the representations of pre-trained individual encoders. Maven (Zhang et al., 2024) is another example, in which simulated and real light curves and spectra of supernovae were used to create a multimodal model for supernova classification and redshift estimation. In addition, AstroM3 (Rizhko & Bloom, 2025) is a model that uses light curves, spectra, and meta-information for stellar variability classification.
@@
The paper is organized as follows: In Section 2, we give general background on self-supervised learning; in Section 3 we present the details of the DESA model; in Section 4 we discuss the dataset and pre-processing steps; in Section 5 we discuss implementation details; in Section 6 we shows the results of our model on various tasks and compare them to various baselines. Finally, Section 7 summarizes the findings and discusses future directions.
2 Self Supervised Learning
2.1 Contrastive learning
@@
Contrastive methods are a subset of self-supervised learning (SSL) techniques that utilize the idea of ‘positive’ and ‘negative’ pairs, aiming to create an embedding space where the positive pairs are close to each other, while the negative pairs are pushed apart. This is performed by using special architectures and contrastive loss functions. Contrastive methods demonstrated great success in the computer vision domain with particularly influential works like SimCLR (Chen et al., 2020) and MoCo (He et al., 2019). While very popular, classical contrastive methods have drawbacks: they require a large batch size to sufficiently represent negative samples at each iteration. Also, the distinction between ‘positive’ and ‘negative’ pairs is a simplified assumption over the data. This is especially important in domains like astrophysics, where the transition between positive and negative pairs is often continuous. In addition, there might be more than one pair of the same object. In particular, in astrophysical data, having more than one spectrum for the same object is very common. Another common issue is the collapse phenomenon that occurs when the model ignores the inputs and creates identical and constant, trivially similar, output vectors. Many works have tried to mitigate those drawbacks in different ways. (Chen & He, 2020) suggested a model called SimSiam with a modification to the classical infoNCE loss used in SimCLR. In SimSiam, the propagation of gradients is performed only through one side of the network. The authors showed experimentally that using this ‘stop gradient’ mechanism allowed the use of the positive pairs only while avoiding the collapse. MoCo uses a dynamical queue and a momentum-based encoder to reduce the need for a large batch size. Similarly, BYOL (Grill et al., 2020) applies a slowly-moving average update to one of the encoders. As of today, contrastive methods are still among the most popular SSL frameworks. As an example, both AstroCLIP (Parker et al., 2024), Maven (Zhang et al., 2024), and AstroM3 (Rizhko & Bloom, 2025) were trained using contrastive methods.
2.2 Regularized methods
@@
A different line of work focuses on feature-level discrimination rather than instance-level discrimination, as in contrastive methods. This idea is motivated by canonical correlation analysis and was suggested as a self-supervised method by Zhang et al. (2021) and Zbontar et al. (2021). The latter was the motivation of the Variance-Invariance-Covariance Regularization (VicReg) architecture (Bardes et al., 2021). VicReg applies three different losses to prevent collapse and maintain alignment, ensuring the variance of the embeddings is sufficiently large (variance term), pushing the covariance between features to be the identity matrix (covariance term), and minimizing the $L_{2}$ distance between pairs of embeddings (invariance term). One advantage of this method is that it does not use negative pairs or asymmetric architectures. In their paper, the authors show that this model outperforms MoCo (He et al., 2019), SimSiam (Chen & He, 2020), and BYOL (Grill et al., 2020) without the need for a large batch size.
@@ 3.1 Hybrid training of individual modalities
Similarly to Parker et al. (2024) and Zhang et al. (2024), we start by training individual modalities separately. However, instead of only self-supervised training, we use a hybrid framework. The hybrid framework adds a supervised head to the self-supervised framework and trains the model with the following loss function -
@@
Figure 1: Upper panel - High-level diagram of the entire model. Lower panel - Detailed diagram of the DualFormer module.
Figure 2: Ablation study results. Upper panel: different attention mechanisms. Lower panel: different uses of $A$ – with and without $A^{T}$ for one of the projections (refer to Section 3.2 for details).
@@ 3.1.1 Spectra encoder
To train the spectra encoder, we start with a CNN encoder followed by two different channels, one for the self-supervised task and one for the supervised task. The self-supervised approach for spectra was chosen to be masked filling. As such, the input is a masked spectrum where $15\%$ of the points are masked and replaced with either a pre-defined value ($80\%$ of the points) or a random value ($20\%$ of the points). The masked spectra are processed through the CNN encoder and then split into two branches: a CNN decoder that produces a filled spectrum, and a branch that consists of a Conformer (Gulati et al., 2020) followed by a 2-layer multi-layer perceptron (MLP) which predicts $T_{\mathrm{eff}}$, $\log g$ and $[\mathrm{Fe}/\mathrm{H}]$. Conformer is a transformer-based architecture with an added convolution between the multi-head self-attention layers. It was shown to be effective in capturing both global and local information, and was used by Pan et al. (2024b) and Kamai & Perets (2025a) in their models. As suggested in Pan et al. (2024b), we modified the standard Conformer by using rotary position embedding (RoPE) (Su et al., 2021) instead of sinusoidal positional encoding. We chose $\mathcal{L}_{\mathrm{ssl}}$ to be the mean squared error (MSE) between the masked and filled spectra. $\mathcal{L}_{\mathrm{sup}}$ is a conformalized quantile regression (CQR) loss (Romano et al., 2019). CQR is a form of quantile regression loss that mitigates one of the biggest problems of standard quantile regression – wrong prediction intervals. In CQR, quantile regression is combined with conformal prediction to create statistically calibrated confidence intervals. This is done by first calculating a ‘conformity score’, which quantifies the error at each interval. This error is then used to calibrate the intervals on the test set.
@@ 3.1.2 Light curve encoder
The light curve encoder was trained in a contrastive-hybrid method. Light curves were augmented into two different views by cropping the light curve at different times. For each view, we calculate the Autocorrelation Function (ACF) and the Fast Fourier Transform (FFT), which potentially encode important information (e.g., identification of periodic and quasi-periodic behaviors arising from stellar rotation of, e.g., eclipsing binaries), and add them as additional channels. Each view is then processed by a backbone that consists of two separate channels: a CNN encoder and a Conformer module. The ACF and FFT channels are sent to the CNN encoder, and the normalized light curve is sent to the Conformer encoder. The embeddings are then combined and processed using a SimSiam method. The SimSiam framework can be described as follows: given the two outputs $e_{1}$ and $e_{2}$ of the backbone network, we apply an MLP network, $f$, such that $z_{1}=f(e_{1})$, $z_{2}=f(e_{2})$. We then project $z_{1}$ onto $z_{2}$ using another MLP $g$. The loss function is the cosine similarity between the projected and the unprojected features:
@@
[…]
@@ 3.2 DualFormer
The next step is to combine the embeddings from the pre-trained individual encoders. Here we are using a novel approach specifically tailored for multi-modality in stellar astrophysics. This is motivated by the observation that the information relationships between light curves and spectra of stars are very different from those found in language and vision modalities. While text describes its corresponding image and vice versa, neither the light curve nor the spectra describes the other. They both partially describe the star, each from a very different perspective. Intuitively, light curves and spectra can be seen as orthogonal views of the star, as the former uses the time domain while the latter uses the frequency domain. Of course, the measurements are not mathematically orthogonal, since the measurements come from different surveys, with different bands and sensitivities, and can be taken at very different times, which complicates the relationship. Nevertheless, this intuition does not exist in other sets of modalities, such as image-text or image-spectra. Moreover, in both text and images (as well as spectra and images such as those in AstroCLIP), the dynamics of the system is not manifested in the data. In constrast, light curves measure time-dependent phenomena by design. This creates a time-dependent information relationship in the case of light curve and spectra alignment.
@@
We train the full model using low-resolution spectra from LAMOST (Zhao et al., 2012; Wang et al., 2022) and light curves from Kepler Mathur et al. (2017).

### v2

@@ Abstract
Stellar astrophysics relies on diverse observational modalities-primarily photometric light curves and spectroscopic data from which fundamental stellar properties are inferred. While machine learning (ML) has advanced analysis within individual modalities, the complementary information encoded across modalities remains largely underexploited. We present DESA (Dual Embedding model for Stellar Astrophysics), a novel multi-modal foundation model that integrates light curves and spectra to learn a unified, physically meaningful latent space for stars. DESA first trains separate modality-specific encoders using a hybrid supervised/self-supervised scheme, and then aligns them through DualFormer, a Transformer-based cross-modal integration module tailored for astrophysical data. DualFormer combines cross- and self-attention, a novel dual-projection alignment loss, and a projection-space eigendecomposition that yields physically structured embeddings.
We demonstrate that DESA significantly outperforms leading unimodal and self-supervised baselines across a range of tasks. In zero- and few-shot settings, DESA’s learned representations recover stellar color–magnitude and Hertzsprung–Russell diagrams with high fidelity ($R^{2}=0.92$ for photometric regressions). In full fine-tuning, DESA achieves state-of-the-art accuracy for binary star detection (AUC = $0.99$, AP = $1.00$) and stellar age prediction (RMSE = $0.94$ Gyr). As a compelling case, DESA naturally separates synchronized binaries from young stars, two populations with nearly identical light curves—purely from their embedded positions in UMAP space, without requiring external kinematic or luminosity information. DESA thus offers a powerful new framework for multimodal, data-driven stellar population analysis, enabling both accurate prediction and novel discovery.
@@ 1 Introduction
Understanding the fundamental properties of stars is central to astrophysics, enabling insights into stellar evolution, galactic structure, and the conditions for planet formation. Traditionally, this effort was done using an analysis of different stellar measurements, such as stellar light curves (photometry) and stellar spectra (spectroscopy). While classical spectra analysis is usually used to predict stellar parameters related to absorption and emission lines such as $T_{\mathrm{eff}},\log g,v\sin i$ and metallicity $\mathrm{Fe}/\mathrm{H}]$ (García Pérez et al., 2016; Wu et al., 2014), classical light curves analysis usually uses spots modulation to find periodicity signatures (Reinhold et al., 2013; McQuillan et al., 2014; Santos et al., 2019; Lu et al., 2020; Santos et al., 2021; Reinhold et al., 2023; Hattori et al., 2025) and magnetic activity (Mathur et al., 2014; Santos et al., 2024). The revolution of deep learning models also affected stellar astrophysics, with various works using data-driven models that learn to predict or classify stellar parameters from observations and simulations. For example, Blancato et al. (2020), Claytor et al. (2024),Kamai & Perets (2025a), and Claytor & Tayar (2025) used deep learning models to predict rotational period from light curves. Pan et al. (2024b) and Zuo et al. (2025) utilized Transformer-based models to predict $\log g$. The use of machine learning in spectral analysis dates back to the pre-deep learning era, with the work of Bailer-Jones (2000). Since then, many works have been done combining deep learning and stellar spectra, with Leung & Bovy (2019), Bai et al. (2020), Olney et al. (2020), Leung & Bovy (2023), Li & Lin (2023), and Koblischke & Bovy (2024) as some examples. For a review on machine learning in astronomy, please refer to Li et al. (2025).
@@
Although the richness of works at the intersection between machine learning and stellar astrophysics, most of them use a single modality, and we refer to them as unimodal models. Multi-modality models, on the contrary, try to combine the information from different modalities of the same object. This approach showed great success in NLP and vision with works like CLIP (Radford et al., 2021) and its variants. Multimodality is of great importance in astrophysics – for example, photometry and spectroscopy reveal partial and very different stellar information; thus, accurate estimation of some stellar properties requires information from both modalities. Despite its importance, there are few works that utilize multi-modal models in astrophysics. AstroCLIP (Parker et al., 2024) is one of such few attempts. The authors used galaxy images and spectra to train a CLIP-like model that aligns the representations of pre-trained individual encoders. Maven (Zhang et al., 2024) is another example, in which simulated and real light curves and spectra of supernovae were used to create a multimodal model for supernova classification and redshift estimation. In addition, AstroM3 (Rizhko & Bloom, 2025) is a model that uses light curves, spectra, and meta-information for stellar variability classification.
@@
The paper is organized as follows: In Section 2, we give general background on multi modal self-supervised learning; in Section 3 we present the details of the DESA model; in Section 4 we discuss the dataset and pre-processing steps; in Section 5 we discuss implementation details; in Section 6 we shows the results of our model on various tasks and compare them to various baselines. Finally, Section 7 summarizes the findings and discusses future directions.
2 Multi-Modal Self-Supervised Learning
Multi-modal self-supervised learning is a subfield of self-supervised learning that focuses on leveraging multiple observations (modalities) of the same object, without relying on explicit labels. For example, this could involve aligning an image with its corresponding description, where the two modalities—vision and language—represent the same object. In the context of astrophysical measurements, multimodality typically refers to different surveys observing the same objects, such as a light curve and a spectrum (as discussed in this work) or a spectrum and image, as in Parker et al. (2024). The key advantage of multimodality over unimodality is the ability to uncover inter-modality relationships that may not be evident within each modality individually. These relationships can enhance performance on complex tasks that require information from both modalities. Given the surge in data volume and diversity across many scientific fields, coupled with a lack of corresponding labels, it is unsurprising that multimodal self-supervised learning has gained significant popularity in various domains (Brown et al., 2025; Cui et al., 2025). The primary challenge in multimodal learning lies in aligning the different modalities. There are numerous approaches to address this challenge—some specifically designed for multimodality, while others are more general self-supervised techniques. In the following, we discuss two of the most widely used self-supervised methods that can also be applied to alignment in multimodal scenarios.
2.1 Contrastive alignment
@@
Contrastive methods are a subset of self-supervised learning (SSL) techniques that utilize the idea of ‘positive’ and ‘negative’ pairs, aiming to create an embedding space where the positive pairs are close to each other, while the negative pairs are pushed apart. This is performed by using special architectures and contrastive loss functions. Contrastive methods demonstrated great success in the computer vision domain with particularly influential works like SimCLR (Chen et al., 2020) and MoCo (He et al., 2019). While very popular, classical contrastive methods have drawbacks: they require a large batch size to sufficiently represent negative samples at each iteration. Also, the distinction between ‘positive’ and ‘negative’ pairs is a simplified assumption over the data. This is especially important in domains like astrophysics, where the transition between positive and negative pairs is often continuous. In addition, there might be more than one pair of the same object. In particular, in astrophysical data, having more than one spectrum for the same object is very common. Another common issue is the collapse phenomenon that occurs when the model ignores the inputs and creates identical and constant, trivially similar output vectors. Many works have tried to mitigate those drawbacks in different ways. Chen & He (2020) suggested a model called SimSiam with a modification to the classical infoNCE loss used in SimCLR. In SimSiam, the propagation of gradients is performed only through one side of the network. The authors showed experimentally that using this ‘stop gradient’ mechanism allowed the use of the positive pairs only while avoiding the collapse. MoCo uses a dynamical queue and a momentum-based encoder to reduce the need for a large batch size. Similarly, BYOL (Grill et al., 2020) applies a slowly-moving average update to one of the encoders. In the context of multimodality, CLIP (Radford et al., 2021) is a pioneering work that uses a contrastive approach to align text and images. As of today, contrastive methods are still among the most popular SSL frameworks for both unimodal and multimodal learning. As an example, both AstroCLIP (Parker et al., 2024), Maven (Zhang et al., 2024), and AstroM3 (Rizhko & Bloom, 2025) were trained using CLIP-like contrastive methods.
2.2 Regularized alignment
@@
A different line of work focuses on feature-level discrimination rather than instance-level discrimination, as in contrastive methods. This idea is motivated by canonical correlation analysis and was suggested as a self-supervised method by Zhang et al. (2021) and Zbontar et al. (2021). The latter was the motivation of the Variance-Invariance-Covariance Regularization (VicReg) architecture (Bardes et al., 2021). VicReg applies three different losses to prevent collapse and maintain alignment, ensuring the variance of the embeddings is sufficiently large (variance term), pushing the covariance between features to be the identity matrix (covariance term), and minimizing the $L_{2}$ distance between pairs of embeddings (invariance term). One advantage of this method is that it does not use negative pairs or asymmetric architectures. In their paper, the authors show that this model outperforms contrastive approaches on both unimodal and multimodal settings.
@@ 3.1 Hybrid training of individual modalities
Similar to Parker et al. (2024) and Zhang et al. (2024), we start by training individual modalities separately. However, instead of only self-supervised training, we use a hybrid framework. The hybrid framework adds a supervised head to the self-supervised framework and trains the model with the following loss function -
@@
Figure 1: Upper panel - High-level diagram of the entire model. Lower panels - Detailed diagrams of the DualFormer module, the spectra encoder, and the light curve encoder.
Figure 2: Ablation study results. Upper panel: different attention mechanisms. Lower panel: different uses of AA – with and without $A^{T}$ for one of the projections (refer to Section 3.2 for details).
@@ 3.1.1 Spectra encoder
To train the spectra encoder, we start with a CNN encoder followed by two different channels, one for the self-supervised task and one for the supervised task. The self-supervised approach for spectra was chosen to be masked filling. As such, the input is a masked spectrum where $15\%$ of the points are masked and replaced with either a pre-defined value ($80\%$ of the points) or a random value ($20\%$ of the points). The masked spectra are processed through the CNN encoder and then split into two branches: a CNN decoder that produces a filled spectrum, and a branch that consists of a Conformer (Gulati et al., 2020) followed by a 2-layer multi-layer perceptron (MLP) which predicts $T_{\mathrm{eff}}$, $\log g$ and $\mathrm{Fe}/\mathrm{H}]$. Conformer is a Transformer-based architecture with an added convolution between the multi-head self-attention layers. It was shown to be effective in capturing both global and local information, and was used by Pan et al. (2024b) and Kamai & Perets (2025a) in their models. As suggested in Pan et al. (2024b), we modified the standard Conformer by using rotary position embedding (RoPE) (Su et al., 2021) instead of sinusoidal positional encoding. We chose $\mathcal{L}_{\mathrm{ssl}}$ to be the mean squared error (MSE) between the masked and filled spectra. $\mathcal{L}_{\mathrm{sup}}$ is a conformalized quantile regression (CQR) loss (Romano et al., 2019). CQR is a form of quantile regression loss that mitigates one of the biggest problems of standard quantile regression – wrong prediction intervals. In CQR, quantile regression is combined with conformal prediction to create statistically calibrated confidence intervals. This is done by first calculating a ‘conformity score’, which quantifies the error at each interval. This error is then used to calibrate the intervals on the test set.
@@ 3.1.2 Light curve encoder
First, we preprocessed the light curves. Preprocessing is crucial because the raw Kepler measurements are noisy and uncalibrated. In addition, light curves themselves may not always be the best input for some tasks. For example, the period of a strictly periodic signal would be much more easily detected in the frequency domain than in the time domain (see for example discussion in Claytor et al. (2022)). One complication with light curve pre-processing is that different methods highlight specific information but usually hinder other types of information. For example, standard z-score normalization enhances periodicity but hinders activity-related patterns. Since we aim to create a general encoder, we decided to stack different normalizations and transformations into multiple input channels. We specify the details of light curve preprocessing in section 4.2.
Since we trained the light curve encoder in a contrastive-hybrid method, each light curve was augmented into two different views by cropping it at different times. For each view, we calculate the Autocorrelation Function (ACF) and the Fast Fourier Transform (FFT), which potentially encode important information (e.g., identification of periodic and quasi-periodic behaviors arising from stellar rotation of, e.g., eclipsing binaries), and add them as additional channels. Each view is then processed by a backbone that consists of two separate channels: a CNN encoder and a Conformer module. The ACF and FFT channels are sent to the CNN encoder, and the normalized light curve is sent to the Conformer encoder. The intuition behind this choice is that, in the frequency domain (ACF and FFT channels), periodicity is detected through peak detection. These tasks are ’short-range’ and are typical for CNNs. Extracting meaningful information from the light curve requires some level of denoising and is more ’long-range’. Therefore, it is better suited for Transformers (Morvan et al., 2022). This architectural choice is similar to the one presented in Kamai & Perets (2025a). They used the same Conformer encoder and an LSTM instead of CNN, and showed that using both the ACF and the normalized light curve together is better than using each one alone for period detection (see Table 2 in their paper). The embeddings are then combined and processed using a SimSiam method. The SimSiam framework can be described as follows: given the two outputs $e_{1}$ and $e_{2}$ of the backbone network, we apply an MLP network, $f$, such that $z_{1}=f(e_{1})$, $z_{2}=f(e_{2})$. We then project $z_{1}$ onto $z_{2}$ using another MLP $g$. The loss function is the cosine similarity between the projected and the unprojected features:
@@
where $\langle\cdot,\cdot\rangle$ and $\lVert\cdot\rVert$ denote, respectively, the standard Euclidean inner product and the $\ell_{2}$ norm. One important aspect of this loss is that the gradients flow only through $z_{1}$ and not through $z_{2}$. In the SimSiam paper, the author refer to this as a ‘stop-gradient’ mechanism. The final self-supervised loss is a symmetrized version of (2)
@@ 3.2 DualFormer
The next step is to combine the embeddings from the pre-trained individual encoders. Here we are using a novel approach specifically tailored for multi-modality in stellar astrophysics. This is motivated by the observation that the information relationships between light curves and spectra of stars are very different from those found in language and vision modalities. While text describes its corresponding image and vice versa, neither the light curve nor the spectra describes the other. They both partially describe the star, each from a very different perspective. Intuitively, light curves and spectra can be seen as orthogonal views of the star, as the former uses the time domain while the latter uses the frequency domain. Of course, the measurements are not mathematically orthogonal, since the measurements come from different surveys, with different bands and sensitivities, and can be taken at very different times, which complicates the relationship. Nevertheless, this intuition does not exist in other sets of modalities, such as image-text or image-spectra. Moreover, in both text and images (as well as spectra and images such as those in AstroCLIP), the dynamics of the system is not manifested in the data. In contrast, light curves measure time-dependent phenomena by design. This creates a time-dependent information relationship in the case of light curve and spectra alignment.
@@
A general concern in multi-modal models is that one modality will obscure the information from another modality. In our case, this might happen because the number of spectra samples used in the pretraining phase is $\sim 30$ times larger than the number of light curves. As a consequence, the spectra encoder is larger than the lightcurve encoder by roughly the same factor (see section 5 for details), and we might have a situation where all the information in the final features comes effectively from the spectra features. First, we need to remember that for basic stellar properties, a spectrum is more informative than a light curve. Indeed, most basic stellar properties (e.g., Teff, FeH, logg, spectral type) are usually detected using spectral lines rather than light curve features. Therefore, it makes sense that there would be more ’spectra-related’ information in the final features. However, we want to make sure that the light curve information is not completely overwhelmed by the spectra information and that using the light curve improves the final performance, especially in tasks that require a sophisticated combination of different stellar parameters rather than basic stellar parameters. In sections 6.3 and 6.4, we test such tasks and see that using the combined model significantly improved the performance compared to spectra-only or lightcurve-only encoders. Another possible test is a sensitivity analysis, not related to a specific task, that identifies which features are most influenced by each modality. For that, we performed the following experiment: We calculated the final features of the test set three times. One time with the full input (light curve and spectra), one time with the light curve replaced with zeros (we call this case spectra-only), and one time with the spectra replaced with zeros (lightcurve-only). We then look at the difference between the full features and spectra-only/lightcurve-only features. The difference is calculated per feature and averaged across all samples. This way, we can trace each individual feature and see for which modality (light curve or spectra only) the difference is smaller. This modality is the more dominant for this specific feature. Figure 16 in Appendix A shows the results of such an experiment. It shows the per-feature difference for both lightcurve-only (gray) and spectra-only (red) features. The blue circles mark the indices where the lightcurve-only features are more dominant. There are $75$ circles which correspond to $\sim 30\%$ of the features, an order of magnitude more than the ratio of lightcurve to spectra samples and number of parameters. This is not negligible and supports the results we find in sections 6.3 and 6.4. Interestingly, we see that while there are small areas with clear dominance of one modality (around indices $150$, for example), in most cases both modalities seem important. This is especially true for the last indices that correspond to the added prior information. We conclude that while there is more spectral information in the final features, the contribution of the light curve encoder is not negligible and especially important in complicated tasks that require combining different information types. We note that while this experiment is convincing, it is not as rigorous as a full ablation study where we completely remove different components and retrain the model. Such an experiment could be explored in future work.
@@
We train the full model using low-resolution spectra from LAMOST (Zhao et al., 2012; Wang et al., 2022) https://doi.org/10.12149/100632 (catalog 10.12149/100632), and light curves from Kepler Mathur et al. (2017) http://dx.doi.org/10.17909/T9488N (catalog 10.17909/T9488N).